OpenStudy (anonymous):

Which fractions are greater than -4/7 ? Choose exactly two answers that are correct. A. -1/2 B. -5/8 C. -2/3 D. - 6/11

OpenStudy (anonymous):

You can divide all of the numerators by the denominators, and then see which answers are greater than -4/7

OpenStudy (anonymous):

-1/2=-0.5 -5/8=-0.625 and so on.

OpenStudy (anonymous):

i don't get it im bad at math

OpenStudy (anonymous):

-2/3=-0.66 -6/11=-0.545 -4/7=-0.571

OpenStudy (anonymous):

\[-\frac{ 4 }{ 7 }< - \frac{ 1 }{ 2 }\] \[-\frac{ 4 }{ 7 }>-\frac{ 5 }{ 8 }\] \[-\frac{ 4 }{ 7 }>-\frac{ 2 }{ 3 }\] \[-\frac{ 4 }{ 7 }<-\frac{ 6 }{ 11 }\] So, which two are correct?

OpenStudy (anonymous):

A and D

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

can u help me with a other one